OpenStudy (anonymous):

1/2 x+1/3 y +6=0, i cant get the slope?

OpenStudy (whpalmer4):

You most certainly can! :-) Rearrange the formula to be in slope-intercept form: \[y = mx + b\]Then you can read off the slope (m) and the y-intercept (b).

OpenStudy (anonymous):

separate y and see what you get

OpenStudy (anonymous):

i used m = y-b/x-0, i get the answer 36/-12

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you have this 1/2 x+1/3 y +6=0 1/3y = -1/2x-6 clear??

OpenStudy (anonymous):

now multiply both sides by 3 you will get y= -3/2x-18 now compare it with y = mx+b you will get the slope

OpenStudy (chmvijay):

slope =m=-3/2

OpenStudy (chmvijay):

3x+2y+36 =0 2y=-3x-36 y=-3/2 X -18

OpenStudy (anonymous):

thanks, guys.

OpenStudy (anonymous):

welcome